Determining Ionic Strength

Unlike strong acids and bases that ionize completely in solution, weak acids or bases partially ionize. The tendency to ionize (i.e., the ionic strength) of a weak acid or base can be quantified in several ways including

Ka or Kb,

pKa or pKb,

and percent ionization.


Part A

Pyridine is a weak base that is used in the manufacture of pesticides and plastic resins. It is also a component of cigarette smoke. Pyridine ionizes in water as follows:

C5H5N+H2O?C5H5NH++OH?

The pKb of pyridine is 8.75. What is the pH of a 0.150M solution of pyridine?

pH =


Part B

Benzoic acid is a weak acid that has antimicrobial properties. Its sodium salt, sodium benzoate, is a preservative found in foods, medications, and personal hygiene products. Benzoic acid ionizes in water:

C6H5COOH?C6H5COO?+H+

The pKa of this reaction is 4.2. In a 0.51Msolution of benzoic acid, what percentage of the molecules are ionized?

percent ionized =
Answer #1

Part-A: Ionization reaction of pyridine in water is

         C5H5N+H2O ---> C5H5NH+ + OH-

Init:    0.150M 0M             0M

Eqm: (0.150-x)M xM             xM

Given pKb of pyridine is 8.75

=> - logKb = 8.75

=> Kb = antilog(-8.75) = 1.78*10-9

Now Kb can also be calculated as

Kb = [C5H5NH+]*[OH-] / [C5H5N] = x*x / (0.150-x) = x2 /(0.150-x)    

= >1.78*10-9 = x2 /(0.150-x)    

Since x << 0.150, (0.150-x) is nearly equals to 0.150

=> x = underroot(0.150*1.78*10-9) = 1.63*10-5 M

=> x = [OH-] = 1.63*10-5 M

=> pOH = - log[OH-] = -log(1.63*10-5) = 4.79

=> pH = 14 – pOH = 14 – 4.79 = 9.21 (answer)

Part-B: Let the degree of ionization of benzoic acid be \alpha

Given [C6H5COOH], C = 0.51M

C6H5COOH --> C6H5COO- + H+

Init: 0.51M 0M 0M

eqm: 0.51(1 - \alpha)M 0.51\alphaM 0.51\alphaM

Given pKa = 4.2

=> - logKa = 4.2

=>Ka = antilog(-4.2) = 6.3*10-5

Now Ka can also be calculated as

Ka = [C6H5COO- ]*[H+] / [C6H5COOH] = ( 0.51\alphaM)*(0.51\alphaM) / 0.51(1 - \alpha)M

=>  6.3*10-5 = ( 0.51\alphaM)*(0.51\alphaM) / 0.51(1 - \alpha)M

Since \alpha<<1 , (1 - \alpha) is nearly equals to 1. Hence

  6.3*10-5 = 0.51\alpha2

=> \alpha = 1.11x10-2

Hence percentage ionisation = \alpha*100 = 1.11 % (answer)